OpenStudy (anonymous):

Medal The graph below shows the growth in the number of spam messages sent daily worldwide. Figures are given in billions. How many times greater is the number of spam messages sent daily in 2004 than the number sent in 2001? Round to the nearest tenth. ? times greater

OpenStudy (anonymous):

Probs gonna need the graph, unless you want me to make an answer up. In that case 12.5

OpenStudy (anonymous):

cool so spam messages sent in 2001 was 4 mesages in 2004 was 8 there both in billions so we dont have to think about that. your question asks how many times greater is 8 than 4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you want a number multiplied by 4 that equals 8 4 x a = 8 what is a?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

well if a is 4 then 4 x 4 = 8? nah, 4 x 4 = 16, so it cant be 4 but its good to guess! so if 4 x 4 is 16, and 16 is twice as big as 8, I suggest trying an answer half as big as 4 i.e. 2

OpenStudy (anonymous):

what is 4 x 2?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

flutter, just realised its 8.8 not 8 in your question. so we have 4 x a = 8.8 divide both sides by 4 a = 8.8/4 a = 2.2 times bigger

OpenStudy (anonymous):

thanks so this is the answer?

OpenStudy (anonymous):

you dont think it is?

OpenStudy (anonymous):

i dont know are u sure its right

OpenStudy (anonymous):

well which part are you not sure about? you have 4 billion in 2001, and 8.8 in 2004 so we have to multiply 4 by something to get 8.8

OpenStudy (anonymous):

so 4 times something equals 8.8 4 x something = 8.8 divide both sides by 4 something = 8.8/4 so what is 8.8/4

OpenStudy (anonymous):

2.2

OpenStudy (anonymous):

awesome, so 8 is 2.2 times bigger than 4 your question wants to know how many times 8 is bigger than 4